### v1

@@ 3.1.1 Multi-Level Encoder
where $D^{+}$ and $D^{-}$ represent the sets of positive and negative triples, respectively, and $y_{\tau}$ is the label (0 or 1) of the triple.
The use of BERT allows the model to leverage contextualized text embeddings, improving the quality of predictions compared to traditional methods that might only use structural information of the KG.
Pretrain-KGE 74 enriches traditional Knowledge Graph Embeddings (KRLs) using BERT pre-trained embeddings, aiming to leverage linguistic context for better entity and relation representation. The method is beneficial in contexts where semantic richness is critical, but it relies heavily on the quality and scope of the training data.
StAR 88 introduces a hybrid model combining textual encoding with graph embedding techniques, addressing the limitations of both approaches by modeling graph elements and their spatial relations. It benefits from enhanced representational efficiency but might struggle with complexity and scalability due to the dual nature of the encoding.
MEM-KGC 63 utilizes a Masked Language Model approach for predicting masked entities in a knowledge graph, enhancing embeddings for better performance on KGC tasks. Its strength lies in handling unseen entities during testing, though it might introduce biases from the pre-training data used in the MLM.
The advantage of this approach lies in leveraging the robust linguistic capabilities of BERT, which can capture complex patterns in text data that are beneficial for knowledge graph completion. However, the reliance on textual representation may limit the model’s ability to capture purely structural relationships inherent in knowledge graphs, potentially leading to performance bottlenecks when dealing with entities that have poor textual descriptions.
3.1.2 Score Function Variation
This category focuses on utilizing pre-trained language models across multiple tasks simultaneously, such as link prediction, entity resolution, and relation classification. Techniques often involve fine-tuning a shared language model on KG-specific data, employing strategies that allow the model to leverage learned representations more effectively. The process of the typical method is depicted in Fig. 5.
Figure 5: An overview of score function variation method.
Papers in this subcategory explore the enhancement of knowledge graph embeddings through multi-task learning frameworks and fine-tuning processes. These approaches often integrate additional tasks like relation prediction and relevance ranking with the primary link prediction task, aiming to create a more holistic learning scenario that captures various aspects of knowledge graphs.
Each KG triplet $(h,r,t)$ is represented by natural language descriptions of the head $h$, relation $r$, and tail $t$. These descriptions are tokenized into sequences $T_{h}$, $T_{r}$, and $T_{t}$. A pre-trained language model (LM) encodes these sequences. The output embeddings for the head, relation, and tail are obtained by mean pooling the LM outputs for respective tokens. The semantic embedding process uses the formula:
$h=\text{MeanPool}(\text{LM}(T_{h}))$ (7)
$r=\text{MeanPool}(\text{LM}(T_{r}))$ (8)
$t=\text{MeanPool}(\text{LM}(T_{t}))$ (9)
where $T_{h},T_{r},T_{t}$ are token sequences for the head, relation, and tail, respectively.
After semantic embedding, a structured loss is optimized to align the embeddings with the KG structure. This loss is designed to make the embedding of a true triplet $(h,r,t)$ such that $h+r$ approximates $t$.
The structure loss is defined by the scoring function:
$f(h,r,t)=b-\frac{1}{2}\|h+r-t\|^{2}$ (10)
where $b$ is a margin parameter. The loss aims to minimize the Euclidean distance between the predicted and actual tail entity in the embedding space.
The model estimates the probability of a triplet being correct based on the structured score function. To efficiently handle large KGs, negative sampling is employed, which involves generating negative triplets by corrupting either the head or the tail entity.
MTL-KGC 66 integrates multiple learning tasks (link prediction, relation prediction, relevance ranking) to enhance KGC using a shared BERT layer. Multi-task learning helps in leveraging linguistic information across different tasks, enhancing overall learning efficiency. However, it may suffer from task interference where one tasks learning negatively impacts another.
SimKGC 86 introduces an efficient contrastive learning strategy with advanced negative sampling for text-based KGC. It focuses on hard negatives to enhance model training, which improves learning efficiency but may need extensive tuning to optimize the negative sampling strategy.
KEPLER 47 fuses the capabilities of PLMs with traditional KE techniques, offering dual optimization on both factual and linguistic fronts. This integration allows for richer representations, though the complexity of joint optimization could lead to challenges in model convergence.
The multi-task learning approach helps in capturing a more comprehensive semantic understanding by tackling multiple aspects of knowledge graph completion simultaneously. This can lead to models that generalize better on unseen data. However, these methods can be computationally expensive and complex to tune, as they require balancing multiple learning objectives which might conflict.
3.1.3 Negative Strategy
As shown in Fig. 6, some integrate textual information encoded by LLMs with structured knowledge graph embeddings. This fusion approach enhances the ability to handle both the semantic nuances captured through text and the explicit relational data represented in knowledge graphs.
Figure 6: An overview of negative strategy method.
Research in this area focuses on hybrid models that combine the strengths of both textual and structural data. Methods typically involve using language models for initial encoding followed by integration with graph embedding techniques, which help in capturing both textual nuances and relational structures within the graph.
Each entity $v$ and relation $e$ in the graph $\mathcal{G}$ is embedded into a $4\mathbf{D}$-dimensional hypercomplex space, capturing both the structural and textual information:
$v=\mathbf{s}_{v}+\mathbf{x}_{v_{1}}\mathbf{i}+\mathbf{y}_{v_{2}}\mathbf{j}+\mathbf{z}_{v_{3}}\mathbf{k}$ (11)
$e=\mathbf{s}_{e}+\mathbf{x}_{e_{1}}\mathbf{i}+\mathbf{y}_{e_{2}}\mathbf{j}+\mathbf{z}_{e_{3}}\mathbf{k}$ (12)
where $\mathbf{s}_{v}$ and $\mathbf{s}_{e}$ represent the structural embeddings, and $\mathbf{x}_{v_{1}},\mathbf{y}_{v_{2}},\mathbf{z}_{v_{3}}$, etc., are embeddings derived from textual information.
The compatibility of a triplet $(v,e,w)$ in $\mathcal{G}$ is scored using a function that measures the alignment of the Dihedron product of the embeddings:
$f(v,e,w)=\mathbf{Re}\left((\overline{v\otimes e})\otimes w\right)$ (13)
where $\otimes$ denotes the Dihedron product, and $\overline{v\otimes e}$ represents the conjugate of the product.
The method stands out by efficiently integrating diverse information sources and leveraging the mathematical properties of hypercomplex spaces to enhance the semantic richness of knowledge graph embeddings.
CoDEx 28 guides traditional KRL models using embeddings derived from LLMs, enhancing the link prediction capabilities. It benefits from the textual richness of LLMs but depends heavily on the initial quality of the language model embeddings.
OpenWorld KGC 72 presents a unified learning framework that combines MLM with multi-task learning to integrate new entities into KGs effectively. It excels in handling unconnected entities but might struggle with data sparsity issues.
Lovelace et al. 61 optimizes PLMs for KGC by decoupling querying and candidate retrieval embeddings, using supervised and unsupervised techniques. It enhances embedding suitability for KGC, but maintaining scalability and efficiency can be challenging.
The integration of textual and structural embeddings allows for a richer representation of knowledge graphs, potentially improving accuracy in tasks such as link prediction and entity resolution. Nevertheless, the complexity of these models can lead to challenges in training stability and increased computational demands.
@@ 3.2 Encoder-Decoder-Based Methods
In this category, LLMs are employed as sequence-to-sequence generators in knowledge graph completion. The models receive textual sequences of query triples and directly generate the text of the tail entity, employing the full generative capabilities of language models to predict missing elements of the graph.
3.2.1 Multi-Level Encoder
Recent works explore advanced contextual embedding strategies and contrastive learning methods to refine knowledge graph completion techniques. The use of rich contextual embeddings from language models combined with innovative training techniques like contrastive loss helps in distinguishing between correct and incorrect triples more effectively. Fig. 7 outlines the steps involved in the conventional method.
Figure 7: An overview of multi-level encoder method.
This subgroup applies contextual embeddings and contrastive learning techniques to knowledge graph completion. Contextual methods leverage language models to generate rich, nuanced embeddings, while contrastive approaches focus on distinguishing between correct and incorrect triples through advanced sampling strategies and loss functions.
For discrimination-based methods, the input sequence for a KG triplet $(v,e,w)$ is prepared as:
$\mathbf{X}_{\text{pair}}=[CLS]\,\mathbf{x}_{v}[SEP]\mathbf{x}_{e}[SEP]$ (14)
$\mathbf{X}_{\text{tail}}=[CLS]\,\mathbf{x}_{w}[SEP]$ (15)
where $\mathbf{x}_{v},\mathbf{x}_{e},\mathbf{x}_{w}$ are the text representations of the head entity, relation, and tail entity, respectively.
For generation-based methods, the scoring function for a KG triplet is defined using the output probabilities of the sequence model:
This represents the likelihood of generating the tail entity sequence given the head entity and relation sequences.
This method represents a significant step forward in KRL research by combining the strengths of PLMs with graph representation learning, thereby improving the effectiveness of knowledge extraction and reasoning tasks.
KG-GPT2 49 employs GPT-2 for KGC, treating triples as sentences for sequence classification. It leverages GPT-2’s contextual understanding but may not always capture the structural nuances of KGs effectively.
GenKGC 40 transitions KGC to a Seq2Seq generation problem using PLMs, focusing on efficient entity generation. It reduces inference times but may sacrifice accuracy for speed.
CD 20 distills context-rich information from LLMs into KGC models, enhancing them with detailed text-based contexts. It improves contextual understanding but requires significant computational resources for the distillation process.
Contextual methods benefit from the deep semantic understanding provided by pre-trained language models, which can significantly enhance the interpretability and performance on complex triples. Contrastive methods improve learning efficiency by focusing on hard negatives. However, these techniques can be sensitive to the choice of negatives and require careful tuning of the contrastive loss parameters.
3.2.2 Soft Prompt
This group utilizes sequence-to-sequence and Transformer architectures to perform tasks such as predicting missing entities in triples or completing parts of the knowledge graph. The flexibility of Seq2Seq models is particularly useful for tasks where structured inputs need to be transformed into textual or other structured outputs. The sequence of actions in the typical method is demonstrated in Fig. 8.
Figure 8: An overview of soft prompt method.
Seq2Seq and Transformer-based models in this category utilize sequence-to-sequence frameworks to generate missing entities or relations in a knowledge graph directly. These models are pre-trained on large text corpora and fine-tuned on knowledge graph data, leveraging their ability to generate coherent and contextually appropriate text sequences.
Each entity, relation, and triplet in the KG is treated as a text sequence. This allows for the use of textual prompts in LLMs for prediction tasks.
$\text{Triple}=(v,e,w)$ (16)
@@ 3.2.2 Soft Prompt
For triple classification task, which involves determining if a given triple $(v,e,w)$ is correct (valid) or not, the prompt example is:

### v2

@@ 3.1.1 Triple Encoding
where $y_{\tau}\in\{0,1\}$ is the label indicating whether the triple is positive or negative, and $D^{+}$ and $D^{-}$ are the sets of positive and negative triples, respectively. The negative triples are generated by corrupting either the head or the tail entity of a positive triple.
MTL-KGC [20] utilizes a multi-task learning framework to enhance knowledge graph completion. It incorporates additional tasks like relation prediction and relevance ranking, significantly improving upon the limitations of prior models such as KG-BERT by learning more relational properties and distinguishing between lexically similar candidates effectively.
K-BERT [21] integrates structured domain knowledge from knowledge graphs directly into the pre-training process. It maintains the original BERT architecture but adds a mechanism to inject relevant triples from a knowledge graph into the training examples. This approach aims to enrich the language model with domain-specific information, addressing BERT’s limitation in handling domain-specific tasks without extensive retraining.
MLMLM [22] introduces a novel method for link prediction by employing Mean Likelihood Masked Language Models to directly generate potential entities. This approach allows for leveraging the knowledge embedded in pre-trained MLMs to enhance the interpretability and scalability of knowledge graphs.
PKGC [23] is a new PLM-based model for KGC that utilizes prompt engineering to enhance knowledge utilization from PLMs. This model adapts PLMs for the KGC task by converting knowledge triples into natural language prompts, thereby facilitating the effective use of PLMs’ latent knowledge.
CSProm-KG [24] innovatively incorporates both structural and textual information into the knowledge graph completion process. This model employs Conditional Soft Prompts that adapt the inputs based on the structure of the knowledge graph, allowing pre-trained language models to efficiently utilize both types of information.
3.1.2 Translation Encoding
Translation encoding methods focus on encoding the head entity and relation together while treating the tail entity separately. This technique relies on optimizing the vector space distances between these entities, facilitating more accurate knowledge representation and reasoning. Fig. 4 depicts a representative model StAR [25] that employs translation encoding.
Figure 4: An overview of translation encoding method.
This model partitions each triple into two asymmetric parts: one combining the head entity and relation, and the other with the tail entity. The StAR model then uses a Siamese-style textual encoder to encode these parts into contextualized representations. The translation function is formulated as follows:
$\mathbf{u}=\text{Pool}(\text{Transformer-Enc}(\mathbf{X}_{(h,r)})),\quad\mathbf{v}=\text{Pool}(\text{Transformer-Enc}(\mathbf{X}_{t}))$ (7)
where $\mathbf{X}_{(h,r)}=[\text{CLS}]\;\mathbf{X}_{h}\;[\text{SEP}]\;\mathbf{X}_{r}\;[\text{SEP}]$ and $\mathbf{X}_{t}=[\text{CLS}]\;\mathbf{X}_{t}\;[\text{SEP}]$ are the input sequences for the head-relation pair and the tail entity, respectively. $\text{Pool}(\cdot)$ denotes the pooling operation to obtain the sequence-level representation from the transformer’s output.
The plausibility score of a triple $(h,r,t)$ is then determined by a combination of a deterministic classifier and a spatial measurement:
$s_{c}=\text{softmax}(\text{MLP}([\mathbf{u};\mathbf{u}\circ\mathbf{v};\mathbf{u}-\mathbf{v};\mathbf{v}])),\quad s_{d}=-\|\mathbf{u}-\mathbf{v}\|_{2}$ (8)
where $[\mathbf{u};\mathbf{u}\circ\mathbf{v};\mathbf{u}-\mathbf{v};\mathbf{v}]$ represents the interactive concatenation of the contextualized embeddings $\mathbf{u}$ and $\mathbf{v}$, and $\text{MLP}(\cdot)$ is a multi-layer perceptron.
The model is trained using both a triple classification objective and a contrastive loss:
$L_{c}=-\frac{1}{|D|}\sum_{t_{p}\in D}\left(\log s_{c}+\sum_{t_{p}^{\prime}\in N(t_{p})}\log(1-s_{c}^{\prime})\right)$ (9)
$L_{d}=\frac{1}{|D|}\sum_{t_{p}\in D}\sum_{t_{p}^{\prime}\in N(t_{p})}\max(0,\lambda-s_{d}+s_{d}^{\prime})$ (10)
where $D$ is the set of positive triples, $N(t_{p})$ denotes the set of negative triples generated by corrupting positive triples, and $\lambda$ is the margin.
The overall training loss is a weighted sum of the two objectives:
$L=L_{c}+\gamma L_{d}$ (11)
where $\gamma$ is a hyperparameter controlling the balance between the two losses.
SimKGC [26] utilizes a bi-encoder architecture to improve contrastive learning efficiency in knowledge graph completion. It integrates three types of negative sampling, in-batch, pre-batch, and self-negatives, enhancing training effectiveness by leveraging larger negative sets and a novel InfoNCE loss function.
LP-BERT [27] employs a multi-task pre-training strategy, enhancing knowledge graph completion by predicting linkage between entities. This method integrates masked language, entity, and relation modeling tasks, significantly improving link prediction. It effectively handles unseen entities and relations by leveraging contextualized entity and relation information from large datasets.
3.1.3 Independent Encoding
Independent encoding methods separately encode each component of a triple: the head entity, relation, and tail entity. This approach allows for flexible and modular representations, which can be beneficial for certain types of knowledge graph applications. The architecture of a typical independent encoding model KEPLER [28] is shown in Fig. 5.
Figure 5: An overview of independent encoding method.
The KEPLER model employs the following methodology for entity descriptions as embeddings:
$h=E_{\langle s\rangle}(\text{texth}),\quad t=E_{\langle s\rangle}(\text{textt}),\quad r=T_{r},$ (12)
where texth and textt are the descriptions for the head and tail entities, respectively, with a special token $\langle s\rangle$ at the beginning. $T\in\mathbb{R}^{|R|\times d}$ is the relation embeddings matrix, and $h$, $t$, and $r$ are the embeddings for the head entity, tail entity, and relation.
The knowledge embedding objective of KEPLER uses negative sampling for optimization:
$L_{KE}=-\log\sigma(\gamma-d_{r}(h,t))-\frac{1}{n}\sum_{i=1}^{n}\log\sigma(d_{r}(h^{\prime}_{i},t^{\prime}_{i})-\gamma),$ (13)
where $(h^{\prime}_{i},r,t^{\prime}_{i})$ are negative samples, $\gamma$ is the margin, $\sigma$ is the sigmoid function, and $d_{r}$ is the scoring function:
$d_{r}(h,t)=\|h+r-t\|_{p},$ (14)
with the norm $p$ set to 1. The negative sampling policy fixes the head entity and randomly samples a tail entity, and vice versa.
The model BERT-ResNet [29] extends the capabilities of encoder-based KRL methods by combining BERT with a deep residual network, improving the handling of sparse connectivity in knowledge graphs. This model leverages the robustness of BERT embeddings and the expressiveness of ResNet’s deep convolutional architecture, significantly enhancing entity ranking performance even with limited training data.
BLP [30] model focuses on inductive link prediction by utilizing textual entity descriptions and pre-trained language models, emphasizing its ability to generalize well across unseen entities. This model outperforms previous approaches by incorporating dynamic graph embeddings, allowing it to effectively adapt to continuously evolving knowledge graphs without retraining.
@@ 3.2 Encoder-Decoder-Based Methods
Encoder-decoder-based methods utilize models such as BART [2] and T5 [31]. These methods are categorized based on the type of input sequence used by the model. Structural Encoding methods (§3.2.1) are similar to Encoder-based methods and use encoder-type triple sequences as input. Fine-tuning methods (§3.2.2), similar to Decoder-based methods, use natural language expressions of triples as input for the decoder.
3.2.1 Structural Encoding
Structural encoding leverages the structure of the triples, feeding them into the encoder as sequences. This method captures the syntactic and relational structure of the triples, enhancing the overall representation learned by the model. Fig. 6 demonstrates the GenKGC [32] architecture that exemplifies structural encoding.
Figure 6: An overview of structural encoding method.
This method converts knowledge graph completion into a sequence-to-sequence generation task. It uses pre-trained language models to generate target entities from input sequences representing head entities and relations. The structural encoding is described as follows:
Given a triple with a missing tail entity $(e_{h},r,?)$, the input sequence is constructed by concatenating the descriptions of the head entity $d_{e_{h}}$ and the relation $d_{r}$:
@@ 3.2.1 Structural Encoding
where $d_{e_{h}}$ and $d_{r}$ are the textual descriptions of the head entity and the relation, respectively. The output sequence corresponds to the target tail entity $d_{e_{t}}$.
The generative process is defined as:
$p_{\theta}(y|x)=\prod_{i=1}^{N}p_{\theta}(y_{i}|y_{<i},x)$ (16)
where $x$ is the input sequence, $y$ is the output sequence, and $N$ is the length of the output sequence. The model is trained using a standard sequence-to-sequence objective function:
$L=-\log p_{\theta}(y|x)$ (17)
During training, GenKGC employs relation-guided demonstration and entity-aware hierarchical decoding to improve representation learning and reduce inference time. Relation-guided demonstration augments the input sequence with examples of triples that share the same relation:
$x=[\text{CLS}]\;\text{demonstration}(r)\;[\text{SEP}]\;d_{e_{h}}\;[\text{SEP}]\;d_{r}\;[\text{SEP}]$ (18)
Entity-aware hierarchical decoding constrains the decoding process by utilizing entity type information, which helps in reducing the search space during generation and improving inference efficiency.
LambdaKG [33] advances the application of PLM in knowledge graph embeddings. Unlike earlier methods that merely transform triples into embeddings, LambdaKG encapsulates both structural and textual information into a unified model architecture. It harnesses sophisticated training techniques like prompt engineering, negative sampling, and relation-aware modeling to enhance the efficiency and accuracy of knowledge graph representations.
3.2.2 Fine-tuning
Fine-tuning methods adapt pre-trained encoder-decoder models using natural language descriptions of triples. This approach tailors the model’s generative capabilities to specific knowledge representation tasks, improving performance on downstream applications. An illustration of a typical fine-tuning model KGT5 [34] is presented in Fig. 7.
Figure 7: An overview of fine-tuning method.
KGT5 fine-tunes a T5 model on both link prediction and question answering tasks. The process involves first pre-training on link prediction and then fine-tuning on QA datasets, balancing between these tasks with a regularization approach.
Given a set of entities $E$ and relations $R$, a knowledge graph $K\subseteq E\times R\times E$ consists of triples $(s,p,o)$. KGT5 treats KG link prediction and QA as sequence-to-sequence tasks. The model uses textual representations of entities and relations and a verbalization scheme to convert link prediction queries into textual queries.
The training process can be described with the following formulas:
$\text{Score}(a)=p_{a}+\alpha\quad\text{if}\;a\in N(e)$ (19)
$\text{Score}(a)=p_{a}\quad\text{otherwise}$ (20)
where $p_{a}$ is the log probability of the predicted entity $a$, $\alpha$ is a constant hyperparameter, and $N(e)$ represents the $n$-hop neighborhood of the entity $e$.
The model’s architecture and training process can significantly reduce the model size while maintaining or improving performance on large KG and KGQA tasks.
KG-S2S [35] is a state-of-the-art Seq2Seq generative framework designed to tackle various challenges in knowledge graph completion (KGC) across different settings without necessitating structural modifications for each graph type. This model uniquely addresses the limitation of previous KGC approaches that are tightly coupled with specific graph structures, which restricts their adaptability to new or evolving KG. By treating all elements of knowledge graphs—entities, relations, and even metadata—as sequences in a unified ‘flat’ text format, KG-S2S avoids the complexity inherent in traditional graph-based methods. This innovation not only simplifies the input data representation but also enhances the model’s flexibility and scalability. It employs advanced fine-tuning techniques on pre-trained language models, integrating novel mechanisms such as entity descriptions and relational soft prompts to enrich the model’s contextual understanding.
3.3 Decoder-Based Methods
Decoder-based methods, utilizing models such as LLaMA [36] and GPT-4 [37], are characterized by the role of the decoder in the representation learning process. Description Generate methods (§3.3.1) create descriptive text for low-resource entities while using Encoder-based or Encoder-Decoder-based methods for other tasks. Prompt Engineering (§3.3.2)treats the decoder as a question-answering tool, using natural language to retrieve triple information and perform downstream tasks. Structural Fine-tuning (§3.3.3) combines structural and textual embeddings, optimizing the decoder’s output for improved knowledge representation.
3.3.1 Description Generate
Description generate methods enhance the representation of low-resource entities by creating descriptive text. This approach supplements the textual information, improving the performance of previous methods and ensuring a richer, more complete entity representation. Fig. 8 illustrates a typical model CD [38] used in description generate method.
Figure 8: An overview of description generate method.
This approach leverages LLM to transform compact structural triplets into context-rich segments. The process involves generating descriptive contexts for knowledge graph completion using prompts. The key steps can be summarized as follows:
Given a triplet $(h,r,t)$ where $h$ is the head entity, $r$ is the relation, and $t$ is the tail entity, the method generates descriptive context $c$ using an LLM prompted with the triplet. The process can be expressed as:
$p_{i}=\text{Template}(h_{i},r_{i},t_{i})$ (21)
$c_{i}=\text{LLM}(p_{i})$ (22)
where $p_{i}$ is the prompt template filled with the triplet components, and $c_{i}$ is the descriptive context generated by the LLM.
Two auxiliary tasks, reconstruction and contextualization, are introduced to train smaller KGC models with these enriched triplets. The reconstruction task aims to restore corrupted descriptive contexts using masked language modeling (MLM), while the contextualization task trains the model to generate descriptive context from the original triplet. The losses for these tasks are defined as:
$L_{\text{rec}}=\frac{1}{N}\sum_{i=1}^{N}\ell(f(\text{MLM}(c_{i})),c_{i})$ (23)
$L_{\text{con}}=\frac{1}{N}\sum_{i=1}^{N}\ell(f(I_{i}),c_{i})$ (24)
where $\ell$ is the cross-entropy loss function, $f$ is the model, $I_{i}$ is the concatenated input of the head, relation, and tail, and $N$ is the number of samples.
The final loss for training the KGC models combines the KGC loss with the auxiliary task losses:
$L_{\text{final}}=L_{\text{kgc}}+\alpha\cdot L_{\text{rec}}+\beta\cdot L_{\text{con}}$ (25)
where $\alpha$ and $\beta$ are hyperparameters that balance the contributions of the auxiliary losses.
The CP-KGC [39] model represents a significant advancement in the realm of text-based KGC. It specifically leverages Constrained-Prompt Knowledge Graph Completion with Large Language Models to refine and enhance textual descriptions within KGC datasets. CP-KGC utilizes simple, carefully designed prompts that help in regenerating or supplementing existing textual descriptions, thereby improving the overall expressiveness and utility of the data. This approach effectively addresses issues like hallucinations in text generation by LLM, ensuring more accurate and contextually relevant outputs.
3.3.2 Prompt Engineering
Prompt engineering leverages the natural language capabilities of decoders, framing knowledge retrieval and representation tasks as question-answering problems. This method utilizes the decoder to extract and learn from knowledge graph triples efficiently. A representative model KG-LLM [40] using prompt engineering is shown in Fig. 9.
Figure 9: An overview of prompt engineering method.
This method transforms triples into natural language prompts and uses LLM to predict the plausibility of the triples or complete missing information. The process can be described as follows:
For a given triple $(h,r,t)$, where $h$ is the head entity, $r$ is the relation, and $t$ is the tail entity, the prompt is constructed to query the LLM. For instance, for triple classification, the prompt might be: